OpenStudy (anonymous):

Challenge 2 Find the volume of infinite Gabriel's horn. Whether the equation converges or diverges.

OpenStudy (amistre64):

the volume tends to be pi ....

OpenStudy (amistre64):

its the surface are thats infinite

OpenStudy (anonymous):

ya the surface area is infinite....but how do you set the equation to find the volume?

OpenStudy (amistre64):

i believe its \[\Large \lim_{t\to\infty}\int_{1}^{t}\pi(\frac1x)^2~dx\]

OpenStudy (amistre64):

a volume of rotation about the x axis, whose radius is defined by 1/x

OpenStudy (anonymous):

the region it bouds is y=1/x, x-axis, line x=1 and x=b

OpenStudy (amistre64):

x=b is fine too, but since its improper, the limit as b to inf

OpenStudy (anonymous):

hmmm...do you know how to solve the rest of part

OpenStudy (amistre64):

well x^-2 integrates up to -1/x -1/b - -1/1 = 1 - 1/b as b to infinity, 1/b to zero left with pi(1)

OpenStudy (anonymous):

oh that seems easy, so the pi is volume right?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

pi units cubed ... whatever the units may be :)

OpenStudy (anonymous):

so 3.141592^3 right

OpenStudy (anonymous):

and does it converges or diverges

OpenStudy (amistre64):

not pi^3 ... just pi units^3 well, since its goes to pi, id say thats a convergence

OpenStudy (anonymous):

ya thats what i was like it is pi then why is it cubed lol...pi cu. units...haha

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

btw i need one more person to verify what u got

OpenStudy (amistre64):

oh im sure you can find one, there are lots of smarter than mes roaming about :)

OpenStudy (anonymous):

i dont think so...there r few calculus guys around...

OpenStudy (anonymous):

haha.

OpenStudy (anonymous):

well i havent worked on this specifically but i agree with @amistre64 work. everything that he did checks out.

OpenStudy (anonymous):

am i suppose to laugh too? XD lol...now solve it

OpenStudy (anonymous):

i agree too but i wanna make sure with the one who knows

OpenStudy (anonymous):

@amistre64 solved it right though

OpenStudy (anonymous):

it seams to be corect